# existence1.bib

### v1

@@ 4.2 Conditions for Existence of Solutions to General non-convex Problems
The result follows by Theorem 4.2.2 (or Theorem 4.2.3) since a polyhedral set is retractive i.e., $\mathcal{R}(X)=X_{\infty}$ (see Example 3.7). ∎∎
@@ 5 Implications for Convex Problems
In this section, we discuss the relationship between the asymptotic cone and the lineality space [bertsekas_nedić_ozdaglar_2013] of a closed convex set. We also discuss properties of $\mathcal{R}(f)$ for a closed, proper function $f$. Then, we apply our main results from the preceding section to a convex minimization problem and show that our results are more general than those prior.
5.1 Asymptotic Cone, Lineality Space, and Cone of Retractive Directions
For a nonempty closed convex set $X\subseteq\mathbb{R}^{n}$, the asymptotic cone $X_{\infty}$ has two simple characterizations, as follows (Proposition 2.1.5 of [auslender_teboulle_2002]):
$X_{\infty}=\{d\in\mathbb{R}^{n}\mid\exists x\in X\hbox{ such that }x+td\in X,\,\forall t\geq 0\},$
$X_{\infty}=\{d\in\mathbb{R}^{n}\mid x+td\in X,\,\forall x\in X,\,\forall t\geq 0\}.$
The cone $X_{\infty}$ is often referred to as a recession cone of $X$ [rockafellar-1970a, bertsekas_nedić_ozdaglar_2013]. The lineality space of a nonempty closed convex set $X$ is the set defined as follows [rockafellar-1970a, bertsekas_nedić_ozdaglar_2013]:
${\rm Lin}(X)=X_{\infty}\cap(-X_{\infty}),$
or equivalently
@@ 5.1 Asymptotic Cone, Lineality Space, and Cone of Retractive Directions
The inclusion can be strict. For example, if $X=\{x\in\mathbb{R}^{n}\mid Ax\leq b\}$ for some matrix $A$ and a vector $b$, then ${\rm Lin}(X)=\{d\mid Ad=0\}$ and $\mathcal{R}(X)=X_{\infty}=\{d\mid Ad\leq 0\}$ by the polyhedrality of $X$.
The constancy space $\mathcal{C}(f)$ (cf. (6)) of a proper closed convex function $f$ satisfies the following relations (see Theorem 2.5.3 in [auslender_teboulle_2002]):
@@
$\mathcal{C}(f)=\{d\in\mathbb{R}^{n}\mid f(x+td)=f(x),\,\forall x\in{\rm dom}f,t\in\mathbb{R}\}.$
For a proper closed convex function $f$, by the definition of the cone $\mathcal{R}(f)$ of retractive directions of $f$, it follows that
@@
5.2 Sufficient Conditions for Existence of Solutions
We next consider the special case of Theorem 4.2.2 as applied to a general convex problem of the form
@@ 5.2 Sufficient Conditions for Existence of Solutions
We modify Theorem 4.2.2 to obtain the following result.
Theorem 5.2.1.
Let $X$ be a closed convex set, and let $f$ and each $g_{j}$ be a proper closed convex function such that $C\cap(\cap_{j=1}^{m}{\rm dom}g_{j})\cap{\rm dom}f\neq\emptyset$. Assume that
@@
Then, the problem (17) has a finite optimal value and its solution set is nonempty. Moreover, if the set $C$ is polyhedral, then the result holds under a weaker condition that
@@
The set $X$ is nonempty, closed, and convex, so by Proposition 2.2.1, we have
@@
By Proposition 2.3.1, we have that $\{x\mid g_{j}(x)\leq 0\}_{\infty}=\mathcal{K}(g_{j})$ for all $j=1,\ldots,m$, thus implying that
@@
Moreover, by Proposition 3.2.2, we have that
@@
The preceding two relations combined with (19) show that the condition $X_{\infty}\cap\mathcal{K}(f)\subseteq\mathcal{R}(X)\cap\mathcal{R}(f)$ of Theorem 4.2.2 is satisfied, and the result follows.
When the set $C$ is polyhedral, the result follows by Corollary 4.2.1. ∎∎
Theorem 5.2.1 is more general than Proposition 6.5.4 in [bertsekas_nedić_ozdaglar_2013], which requires that the problem (17) has a finite optimal value and that
@@
The preceding condition implies that the condition (19) holds since ${\rm Lin}(C)\subseteq\mathcal{R}(C)$ and the analogous relation holds for the constancy space and the cone of retractive directions for the functions $f$ and $g_{j}$.
For the case of a polyhedral set $C$, Theorem 5.2.1 is more general than Proposition 6.5.5 in [bertsekas_nedić_ozdaglar_2013], which requires that the optimal value of the problem (17) is finite and that a stronger condition than (20) holds, namely that
$C_{\infty}\cap\left(\cap_{j=1}^{m}\mathcal{K}(g_{j})\right)\cap\mathcal{K}(f)\subseteq\left(\cap_{j=1}^{m}\mathcal{C}(g_{j})\right)\cap\mathcal{C}(f).$
Finally, when the set $C$ is polyhedral and $g_{j}\equiv 0$ for all $j$, the condition (20) reduces to
@@
In this case, Theorem 5.2.1 provides a weaker sufficient condition than that of Theorem 27.3 in [rockafellar-1970a] requiring that $C_{\infty}\cap\mathcal{K}(f)\subseteq\mathcal{C}(f)$.
6 Implications for Non-convex Problems
In this section, we consider the implications of our main results for several types of non-convex problems for the case where the constraint set $X$ is generic and the case when $X$ is given by non-convex functional inequalities.
@@ 6.1 Generic Constraint Set
We consider the problem ($P$) for the case when $f$ is convex, for which we have a special case of Theorem 4.2.2.
Theorem 6.1.1.
Let $X$ be a nonempty closed set and $f$ be a proper closed convex function with $X\cap{\rm dom}f\neq\emptyset$. Assume that
@@
Then, the problem ($P$) has a finite optimal value and a solution exists.
@@
A convex function exhibits asymptotically bounded decay with respect to $g(x)=\|x\|$ as seen in Example 3.2. Thus, the result follows from Theorem 4.2.2 and the fact that $\mathcal{C}(f)\subseteq\mathcal{R}(f)$ when $f$ is convex. ∎∎
To the best of our knowledge the result of Theorem 6.1.1 is new. An existing result that considers convex objective and a non-convex constraint set is Proposition 12 in [bertsekas_tseng_2006], which relies on the stringent assumption that:
  1. (A1)
    Every nonzero direction $d\in X_{\infty}$ is retractive and, for all $x\in X$, there exists an $\bar{\alpha}\geq 0$ such that $x+\alpha d\in X$ for all $\alpha\geq\bar{\alpha}.$
In the following example Assumption (A1) fails to hold, so Proposition 12 in [bertsekas_tseng_2006] cannot be applied to assert the existence of solutions. However, Theorem 6.1.1 can be used.
@@
Consider minimizing a proper closed convex function $f$ on the set $X$ given by $X=\{x\in\mathbb{R}^{2}\mid x_{2}\leq x_{1}^{2}\}$ (cf. Figure 2.2). The complement of $X$ is open and convex. Hence, by Proposition 4 of [bertsekas_tseng_2006], we have that $\mathcal{R}(X)=X_{\infty}$. However, the set $X$ does not satisfy Assumption (A1) since, for the direction $d=(0,1)\in X_{\infty}$ and any $x$ that lies on the boundary of $X$ (i.e., $x_{1}^{2}=x_{2}$), it is not the case that $x+\alpha d\in X$ for any $\alpha>0$. Thus, Proposition 12 of [bertsekas_tseng_2006] cannot be used to claim the existence of solutions in this case. However, if $\mathcal{K}(f)={\rm Lin}(f)$ (such as, for example, when $f(x_{1},x_{2})=|x_{1}|)$ then by Theorem 6.1.1, the problem $\inf_{x\in X}f(x)$ has a solution. ∎
Now, we consider the problem $\inf_{x\in X}f(x)$ with a non-convex set $X$ and a non-convex function $f$ and compare our Theorem 4.2.2 with Proposition 13 in [bertsekas_tseng_2006]. We first recast the assumptions of Proposition 13 in [bertsekas_tseng_2006] and then prove that this proposition is a special case of Theorem 4.2.2.
The problem that [bertsekas_tseng_2006] considers is exactly ($P$) under Assumption (A1) and an additional assumption for a sequence of sets $\{S_{k}\}$, such that $S_{k}=X\cap L_{\gamma_{k}}(f)$ where $\{\gamma_{k}\}\subseteq\mathbb{R}$ is a decreasing scalar sequence, which requires that
  • (A2)
    For every asymptotic direction [^1] A (nonzero) direction $d$ such that for some unbounded sequence $\{x_{k}\}$, with $x_{k}\in S_{k}$ for all $k$, we have $\lim_{k\to\infty}x_{k}/\|x_{k}\|=d/\|d\|$. $d$ of $\{S_{k}\}$ and for each $x\in X$,
    either  $\lim_{\alpha\to\infty}f(x+\alpha d)=-\infty$  or  $f(x-d)\leq f(x).$
The statement of Proposition 13 in [bertsekas_tseng_2006] is as follows: Suppose that Assumptions (A1) and (A2) hold for problem ($P$), where the set $X$ is closed and the function $f$ is proper and closed with $X\cap{\rm dom}f\neq\emptyset$. Then, the objective function $f$ attains a minimum over $X$ if and only if the optimal value $f^{*}=\inf_{x\in X}f(x)$ is finite.
We have already discussed the limitations of Assumption (A1) in Example 6.1 that also apply to Proposition 13 in [bertsekas_tseng_2006]. Now we show that Assumptions (A1) and (A2) imply the assumptions of Theorem 4.2.2.
Proposition 6.1.1.
Let $X$ be a closed set and $f$ be a proper closed function with $X\cap{\rm dom}f\neq\emptyset$. Suppose that problem ($P$) satisfies Assumptions (A1) and (A2). If the optimal value $f^{*}=\inf_{x\in X}f(x)$ is finite, then $f$ exhibits asymptotically bounded decay with respect to $g(x)=\|x\|$ on the set $X$. Furthermore, the following relation holds
Proof.
Since $f$ is assumed to have a finite minimum on $X$, by Example 3.1 we have that $f$ exhibits asymptotically bounded decay with respect to the function $g(x)=\|x\|$. Since $\{\gamma_{k}\}$ is a decreasing scalar sequence, we have that
$S_{k}=X\cap L_{\gamma_{k}}(f)\subseteq X\cap L_{\gamma_{1}}(f)=S_{1}\qquad\hbox{for all }k\geq 1.$
Therefore, we have that $d\in(X\cap L_{\gamma_{1}}(f))_{\infty}$. By Proposition 2.2.1, the retractive directions of the intersection set is contained in the intersection of retractive directions of the sets, so we have for any asymptotic direction $d$ of $\{S_{k}\}$,
$d\in(X\cap L_{\gamma_{1}}(f))_{\infty}\subseteq X_{\infty}\cap(L_{\gamma_{1}}(f))_{\infty}.$
Moreover, by Proposition 2.3.1 it holds that $(L_{\gamma_{1}}(f))_{\infty}\subseteq\mathcal{K}(f)$. Thus, for any asymptotic direction $d$ of $\{S_{k}\}$, we have
$d\in X_{\infty}\cap\mathcal{K}(f)=\mathcal{R}(X)\cap\mathcal{K}(f),$ (21)
where the equality holds since $X$ is a retractive set by Assumption (A1).
Next, by Assumption (A2), for every $x\in X$, either $\lim_{\alpha\to\infty}f(x+\alpha d)=-\infty$ or $f(x-d)\leq f(x)$. Suppose that the former case holds for some $x\in X$. Then, as $\alpha\to\infty$, we have that $f(x+\alpha d)\to-\infty$ but this is a contradiction since, by Assumption (A1), $x+\alpha d\in X$ for all $\alpha\geq\bar{\alpha}$ and $f$ was assumed to have a finite minimum on $X$. Hence, we must have
$f(x-d)\leq f(x)\qquad\hbox{ for all $x\in X$}.$ (22)
Let $\{x_{k}\}$ be a sequence associated with an asymptotic direction $d$ of $\{S_{k}\}$ (i.e., $x_{k}\in S_{k}$ for all $k$ with $x_{k}/\|x_{k}\|^{-1}\to d/\|d\|$). Note that if $x_{k}/\|x_{k}\|\to d/\|d\|$, then we also have for any $\rho>0$,
$\lim_{k\to\infty}\frac{x_{k}}{\|x_{k}\|}=\frac{d}{\|d\|}=\frac{\rho d}{\|\rho d\|}.$
Therefore, $\rho d$ is also an asymptotic direction of $\{S_{k}\}$ and relation (22) must also hold for any asymptotic direction $d$ of $\{S_{k}\}$ and for any $\rho$,
$f(x-\rho d)\leq f(x)\qquad\hbox{ for all $x\in X$}.$
In particular, it holds for every point $x_{k}$ of the sequence $\{x_{k}\}$, thus implying that $d$ is a retractive direction of the function $f$ according to Definition 3.3.1. Hence, every asymptotic direction $d$ of $\{S_{k}\}$ also lies in the set $\mathcal{R}(f)$. This and relation (21), imply that $d\in\mathcal{R}(X)\cap\mathcal{R}(f)$. ∎∎

### v3

@@ 4.1 Sufficient Conditions for Existence of Solutions
The result follows by Theorem 4.2 since a polyhedral set is retractive i.e., $\mathcal{R}(X)=X_{\infty}$ (see Example 3.7). ∎
Remark 1.
The condition $X_{\infty}\cap\mathcal{K}(f)\subseteq\mathcal{R}(X)\cap\mathcal{R}(f)$ of Theorem 4.2 implies that we must have
$X_{\infty}\cap\mathcal{K}(f)=\mathcal{R}(X)\cap\mathcal{R}(f),$
since $\mathcal{R}(X)\cap\mathcal{R}(f)\subseteq X_{\infty}\cap\mathcal{K}(f)$ always holds by the definition of the retractive cones.
We next provide a result for the special case when the conditions of Theorem 4.2 hold with $g(x)=\|x\|^{p}$ for some $p\geq 0$. In this case, we can have $p=0$ so that $g$ need not be coercive, and the Lipschitz continuity of $\nabla g$ is not needed.
Theorem 4.3.
Let the set $X\subseteq\mathbb{R}^{n}$ be closed, and let the function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{+\infty\}~{}$be proper and closed with $X\cap{\rm dom}f\neq\emptyset$. Assume that $f$ exhibits asymptotically bounded decay with respect to $g(x)=\|x\|^{p}$ on $X$ for some $p\geq 0$, and assume that
Then, the problem ($P$) has a finite optimal value $f^{*}$ and a nonempty solution set $X^{*}$.
Proof.
The proof is along the same lines as that of Theorem 4.2, which we follow and point to simplifications due to the choice $g(x)=\|x\|^{p}$. By our assumption that $f$ exhibits asymptotically bounded decay with respect to $g(x)=\|x\|^{p}$ on $X$ for some $p\geq 0$, we have that
If $c=+\infty$, then $f$ is coercive on $X$ so by Proposition 2.4, the problem ($P$) has an optimal solution. So, consider the case $c\in\mathbb{R}$.
implying that
Thus, for any $\rho>0$, the function $f(x)+r\|x\|^{p+1}$ is coercive, so by Proposition 2.4, the regularized problem $\inf_{x\in X}\{f(x)+r\|x\|^{p+1}\}$ has a solution for every $r>0$.
(Step 2: A sequence of solutions to regularized problems is bounded.) This part is the same as Step 2 of the proof of Theorem 4.2, where $g^{2}(x)$ is replaced with $\|x\|^{p+1}$ up to and including relation (12), which in this case reduces to
$\|x_{k}^{*}\|^{p+1}\leq\|x_{k}^{*}-\rho d\|^{p+1}\qquad\hbox{for $k\geq K$}.$
Therefore,
$\|x_{k}^{*}\|^{2}\leq\|x_{k}^{*}-\rho d\|^{2}\qquad\hbox{for all $k\geq K$}.$
On the other hand, since $\lim_{k\to\infty}x_{k}^{*}\cdot\|x_{k}^{*}\|^{-1}=d$ and $\|d\|=1$, we obtain that
$1\leq 0$, which is a contradiction. Hence, it must be that $\{x_{k}^{*}\}$ is bounded.
(Step 3: Any accumulation point of sequence $\{x_{k}^{*}\}$ is a solution to ($P$).) The proof is the same as Step 3 of the proof of Theorem 4.2, where $g^{2}(x)$ is replaced with $\|x\|^{p+1}$. ∎
Next, we provide an example showing that, if the condition $X_{\infty}\cap\mathcal{K}(F)\subseteq\mathcal{R}(X)\cap\mathcal{R}(f)$ is violated, the problem may not have a finite optimal value; hence no solution.
Example 4.1.
Consider the problem of minimizing a convex scalar function $f(x)=-\sqrt{x}$ over its domain $X=\{x\mid x\geq 0\}$. The optimal value is $f^{*}=-\infty$ and there is no solution. Since the function is convex, by Example 3.2, it exhibits asymptotically bounded decay with respect to $g(x)=|x|$ on $X$. The set $X$ is a closed convex cone, and we have $X_{\infty}=X$. The asymptotic cone of $f$ coincides with $X$ i.e., $\mathcal{K}(f)=X$, implying that $X_{\infty}\cap\mathcal{K}(f)=X.$ The cone of retractive directions of $X$ coincides with $X$, since $X$ is a polyhedral set (see Example 3.7). The cone $\mathcal{R}(f)$ of retractive directions of $f$ contains only the zero vector. To see this note that for $x_{k}=\lambda k$, with $\lambda>0$, and $t_{k}=k$ for all $k\geq 1$, we have that $\{x_{k}\}$ converges in the direction $d=\lambda$. However, for any $\rho\in(0,1)$ and any $k\geq 1$,
$f(x_{k}-\rho\lambda)=-\sqrt{k-\rho\lambda}>-\sqrt{k}=f(x_{k}).$
Hence, we have $\mathcal{R}(f)=\{0\}$, and $\mathcal{R}(X)\cap\mathcal{R}(f)=\{0\},$ thus implying that the condition $X_{\infty}\cap\mathcal{K}(f)\subseteq\mathcal{R}(X)\cap\mathcal{R}(f)$ of Theorem 4.3 is violated.
Example 4.2.
Consider the problem of minimizing a convex scalar function $f(x)=e^{x}$ over its domain $X=\mathbb{R}$. The optimal value is $f^{*}=0$ and there is no solution. The function is convex, so by Example 3.2, it exhibits asymptotically bounded decay with respect to $g(x)=|x|$ on $X$. We have $X_{\infty}=X$ and $\mathcal{R}(X)=X$. The asymptotic cone of $f$ coincides is given by
$\mathcal{K}(f)=\{x\mid x\leq 0\},$
$f(x_{k}+\rho\lambda)=e^{k+\rho\lambda}>e^{k}=f(x_{k}).$
Hence, we have $\mathcal{R}(f)=\{0\}$ and $\mathcal{R}(X)\cap\mathcal{R}(f)=\{0\},$ implying that the condition $X_{\infty}\cap\mathcal{K}(f)\subseteq\mathcal{R}(X)\cap\mathcal{R}(f)$ of Theorem 4.3 is violated.
As the preceding examples show, when the condition on the retractive cones of Theorem 4.3 are violated, the problem ($P$) may not have finite optimal value $f^{*}$, or it may have finite $f^{*}$ but not a solution. This leads us to conjecture that the sufficient conditions of Theorem 4.3, and Theorems 4.14.2 might also be necessary. It turns out that it is indeed the case, as seen in the next section.
4.2 Conditions of Theorems 4.14.3 are also Necessary
To show that the conditions of Theorems 4.14.3 are also necessary for the existence of solutions to problem ($P$), we use Corollary 4.1 in [17], which provides necessary and sufficient conditions for solution existence to problems in a Banach space. We restate that corollary for the case of $\mathbb{R}^{n}$.
Proposition 4.4 (Corollary 4.1 in [17]).
Let the set $X\subseteq\mathbb{R}^{n}$ be closed, and let the function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{+\infty\}~{}$be proper and closed. Suppose that $h$ is a coercive function and that $X\cap{\rm dom}f\cap{\rm dom}h\neq\emptyset$. Then, the following two conditions are necessary and sufficient in order that the function $f$ attains its infimum on $X$:
  • a)
    For any (nonzero) $d\in X_{\infty}$, for any sequence $\{x_{k}\}\subset X$, such that $f(x_{k})\to\inf_{x\in X}f(x)$ and $x_{k}/\|x_{k}\|\to d$, for all $k$ large enough there exists $y_{k}\in X$ such that
    $f(y_{k})\leq f(x_{k}),\qquad h(y_{k})<h(x_{k});$
  • b)
    The function $f$ is bounded below on the set $X$.
We note that when $X_{\infty}=\{0\}$ then the condition (a) is vacuous and, thus, satisfied by default. Using Proposition 4.4, we obtain the following result.
Theorem 4.5.
The conditions of Theorems 4.14.3 are also necessary for the existence of solutions to problem ($P$).
Proof.
To show the necessity, we prove that the conditions of each of Theorems 4.14.3 imply that the conditions of Proposition 4.4 are satisfied. Since the conditions of Proposition 4.4 are necessary, the conditions of Theorems 4.14.3 are also necessary.
We note that under the conditions of Theorems 4.14.3, we have that $f^{*}=\inf_{x\in X}f(x)$ is finite. Hence, $f$ is bounded below on $X$, and the condition (b) of Proposition 4.4 is satisfied. It remains to show that the condition (a) of Proposition 4.4 is satisfied. For this, we let $d\in X_{\infty}$ and $\{x_{k}\}\subset X$ be any sequence such that $f(x_{k})\to\inf_{x\in X}f(x)$ and $x_{k}/\|x_{k}\|\to d$. Then, $d\in X_{\infty}$ and, since $\{x_{k}\}\subset X$,
$f_{\infty}(d)\leq\liminf_{k\to\infty}\frac{f(x_{k})}{\|x_{k}\|}\leq 0,$
where the last inequality follows from either $\inf_{x\in X}f(x)$ being finite or equal to $\infty$- ∞. Hence, $d\in\mathcal{K}(f)$, and consequently
$d\in X_{\infty}\cap\mathcal{K}(f).$
Now, we consider each of Theorems 4.14.3 separately.
Case of Theorem 4.1: The condition $X_{\infty}\cap\mathcal{K}(f)=\{0\}$ implies that we must have $d=0$. Thus, there is no nonzero vector satisfying the condition (a) of Proposition 4.4. Hence, the condition is satisfied by default.
Case of Theorems 4.24.3: By the condition $X_{\infty}\cap\mathcal{K}(f)\subseteq\mathcal{R}(X)\cap\mathcal{R}(f)$ of Theorems 4.24.3, and the definitions of cones of retractive directions for $X$ and $f$, for the sequence $\{x_{k}\}\subset X$, with $x_{k}/\|x_{k}\|\to d$, and for any $\rho>0$, there exists $K$ (depending on $\rho$) such that
$x_{k}-\rho d\in X,\qquad f(x_{k}-\rho d)\leq f(x_{k})\qquad\hbox{for all }k\geq K.$
Hence, the condition $f(y_{k})\leq f(x_{k})$ in (a) of Proposition 4.4 is satisfied with $y_{k}=x_{k}-\rho d\in X$ (for an arbitrary $\rho>0$.)
We next show that the condition $h(y_{k})<h(x_{k})$ is also satisfied with $y_{k}=x_{k}-\rho d$ for a larger index $K$, for the function $h(x)=\|x\|^{2}$.
Since $\|d\|=1$, we have
$\|x_{k}-\rho d\|^{2}=\|x_{k}\|^{2}+\rho^{2}-2\rho\langle x_{k},d\rangle=\|x_{k}\|^{2}+\rho\|x_{k}\|\left(\frac{\rho}{\|x_{k}\|}-2\left\langle\frac{x_{k}}{\|x_{k}\|},d\right\rangle\right).$
Since $\|x_{k}\|\to\infty$ and $x_{k}/\|x_{k}\|\to d$, it follows that
$\lim_{t\to\infty}\left(\frac{\rho}{\|x_{k}\|}-2\left\langle\frac{x_{k}}{\|x_{k}\|},d\right\rangle\right)=-2.$
Thus, there exists an index $K_{1}\geq K$ and an $\epsilon\in(0,2)$ so that
$\frac{\rho}{\|x_{k}\|}-2\left\langle\frac{x_{k}}{\|x_{k}\|},d\right\rangle\leq\epsilon-2<0\qquad\hbox{for all }k\geq K_{1}.$
Hence, for $k\geq K_{1}$ and $y_{k}=x_{k}-\rho d$, we have $\|y_{k}\|^{2}<\|x_{k}\|^{2}$ for all $k\geq K_{1}$. Thus, the condition $h(y_{k})<h(x_{k})$ in (a) of Proposition 4.4 is satisfied with $h(x)=\|x\|^{2}$. ∎
By Theorem 4.5 and Corollary 4.1, the following result holds when $X$ is polyhedral.
Corollary 4.2.
The assumptions of Corollary 4.1 are necessary for the problem ($P$), with a polyhedral set $X$, to have a solution.
@@ 5 Implications for Convex Problems
In this section, we discuss the relationship between the asymptotic cone and the lineality space [6] of a closed convex set.
Then, we apply Theorem 4.2 to provide necessary and sufficient conditions for existence of solutions to a convex minimization problem.
For a nonempty closed convex set $X\subseteq\mathbb{R}^{n}$, the asymptotic cone $X_{\infty}$ has a simple characterization (Proposition 2.1.5 of [2])
$X_{\infty}=\{d\in\mathbb{R}^{n}\mid\exists x\in X\hbox{ such that }x+td\in X,\,\forall t\geq 0\}.$
The asymptotic cone $X_{\infty}$ is often referred to as a recession cone of $X$ [18, 6]. The lineality space of a nonempty closed convex set $X$ is the set defined as ${\rm Lin}(X)=X_{\infty}\cap(-X_{\infty})$  [18, 6], and satisfies
@@
The inclusion can be strict. For example, if $X=\{x\in\mathbb{R}^{n}\mid Ax\leq b\}$ for some matrix $A$ and a vector $b$, then ${\rm Lin}(X)=\{d\mid Ad=0\}$ and $\mathcal{R}(X)=X_{\infty}=\{d\mid Ad\leq 0\}$ since $X$ is polyhedral.
The constancy space $\mathcal{C}(f)$ (cf. (6)) of a proper closed convex function $f$ satisfies the following relations (see Theorem 2.5.3 in [2]):
@@
or equivalently $\mathcal{C}(f)=\{d\in\mathbb{R}^{n}\mid f(x+td)=f(x),\,\forall x\in{\rm dom}f,t\in\mathbb{R}\}.$ For a proper closed convex function $f$, by the definition of the cone $\mathcal{R}(f)$ of retractive directions of $f$, it follows that
@@
5.1 Necessary and Sufficient Conditions for Solution Existence
We next consider a general convex problem of the form
@@ 5.1 Necessary and Sufficient Conditions for Solution Existence
We have the following result.
Theorem 5.1.
Let $X$ be a closed convex set, and let $f$ and each $g_{j}$ be a proper closed convex function such that $C\cap(\cap_{j=1}^{m}{\rm dom}g_{j})\cap{\rm dom}f\neq\emptyset$. Then, the problem (15) has a finite optimal value and its solution set is nonempty if and only if
@@
Moreover, if the set $C$ is polyhedral, then the problem (15) has a finite optimal value and its solution set set is nonempty if and only if
@@
The set $X$ is nonempty, closed, and convex, so by Proposition 2.1, we have
@@
By Proposition 2.3, we have that $\{x\mid g_{j}(x)\leq 0\}_{\infty}=\mathcal{K}(g_{j})$ for all $j=1,\ldots,m$, thus implying that
@@
Moreover, by Proposition 3.2, we have that
@@
The preceding two relations combined with (17) show that the condition $X_{\infty}\cap\mathcal{K}(f)\subseteq\mathcal{R}(X)\cap\mathcal{R}(f)$ of Theorem 4.3 is satisfied, and the sufficiency part follows by Theorem 4.3 and the necessary part by Theorem 4.5.
When $C$ is polyhedral, the result follows by Corollary 4.2 and Theorem 4.5. ∎
Theorem 5.1 is more general than Proposition 6.5.4 in [6], which provides only sufficient conditions for the existence of solutions, under more stringent assumptions than those of Theorem 5.1 requiring that problem (15) has a finite optimal value and
@@
The preceding condition implies that the condition (17) holds since ${\rm Lin}(C)\subseteq\mathcal{R}(C)$, and the analogous relation holds for the constancy space and the cone of retractive directions for the functions $f$ and $g_{j}$. Similarly, for the case of a polyhedral set $C$, Theorem 5.1 is more general than Proposition 6.5.5 in [6].
Moreover, when the set $C$ is polyhedral and $g_{j}\equiv 0$ for all $j$, the condition (18) reduces to
@@
In this case, Theorem 5.1 is more general than Theorem 27.3 in [18], which provides only sufficient conditions using a more restrictive condition that $C_{\infty}\cap\mathcal{K}(f)\subseteq\mathcal{C}(f)$.
6 Implications for Nonconvex Problems
In this section, we consider the implications of our main results for several types of nonconvex problems for the cases where the constraint set $X$ is generic and given by nonconvex functional inequalities (Sections 6.1 and 6.2, respectively).
@@ 6.1 Generic Constraint Set
We consider the problem ($P$) for the case when $f$ is convex while the set $X$ need not be convex, for which we have by Theorem 4.3 and Theorem 4.5.
Theorem 6.1.
Let $X$ be a nonempty closed set and $f$ be a proper closed convex function with $X\cap{\rm dom}f\neq\emptyset$. Then, the problem ($P$) has a finite optimal value and a solution exists if and only if
@@
A convex function exhibits asymptotically bounded decay with respect to $g(x)=\|x\|$ (Example 3.2), and we have $\mathcal{C}(f)\subseteq\mathcal{R}(f)$ when $f$ is proper closed convex function. The sufficiency part follows by Theorem 4.3, while the necessary part follows from Theorem 4.5. ∎
To the best of our knowledge the result of Theorem 6.1 is new. An existing result that considers convex objective and a nonconvex constraint set is Proposition 12 in [7], which relies on the stringent assumption that:
(A1) Every nonzero direction $d\in X_{\infty}$ is retractive and, for all $x\in X$, there exists an $\bar{\alpha}\geq 0$ such that $x+\alpha d\in X$ for all $\alpha\geq\bar{\alpha}.$
In the following example assumption (A1) fails to hold. Thus, Proposition 12 in [7] cannot be applied to assert the existence of solutions, while Theorem 6.1 can be used.
@@
Consider minimizing a proper closed convex function $f$ on the set $X$ given by $X=\{x\in\mathbb{R}^{2}\mid x_{2}\leq x_{1}^{2}\}$ (cf. Figure 2.2). The complement of $X$ is open and convex. Hence, by Proposition 4 of [7], we have that $\mathcal{R}(X)=X_{\infty}$. However, the set $X$ does not satisfy assumption (A1) since, for the direction $d=(0,1)\in X_{\infty}$ and any $x$ that lies on the boundary of $X$ (i.e., $x_{1}^{2}=x_{2}$), it is not the case that $x+\alpha d\in X$ for any $\alpha>0$. Thus, Proposition 12 of [7] cannot be used to claim the existence of solutions in this case. However, if $\mathcal{K}(f)={\rm Lin}(f)$ (such as, for example, when $f(x_{1},x_{2})=|x_{1}|)$ then by Theorem 6.1, the problem $\inf_{x\in X}f(x)$ has a solution. ∎